A statistical program is recommended.

Consider the following data for two variables, x and y.

x 9 32 18 15 26
y 11 19 20 16 22

(a) Develop an estimated regression equation for the data of the form ŷ = b0 + b1x. (Round b0 to two decimal places and b1 to three decimal places.)

ŷ =

Comment on the adequacy of this equation for predicting y. (Use α = 0.05.)

The high p-value and low coefficient of determination indicate that the equation is inadequate.

The high p-value and high coefficient of determination indicate that the equation is adequate.  

The low p-value and low coefficient of determination indicate that the equation is inadequate.

The low p-value and high coefficient of determination indicate that the equation is adequate.

(b)Develop an estimated regression equation for the data of the form ŷ = b0 + b1x + b2x2. (Round b0 to two decimal places and b1 to three decimal places and b2 to four decimal places.)

ŷ =

Comment on the adequacy of this equation for predicting y. (Use α = 0.05.)

The high p-value and low coefficient of determination indicate that the equation is inadequate.

The high p-value and high coefficient of determination indicate that the equation is adequate.    

The low p-value and low coefficient of determination indicate that the equation is inadequate.

The low p-value and high coefficient of determination indicate that the equation is adequate.

(c) Use the model from part (b) to predict the value of y when x = 20.  (Round your answer to two decimal places.)
Answer #1

applying regression from excel:data -data analysis:

a)

y^ =10.39+0.361x

The high p-value and low coefficient of determination indicate that the equation is inadequate.

b)

y^ =-5.33+2.155x-0.0433x2 \

The low p-value and high coefficient of determination indicate that the equation is adequate.

c)

predicted value =-5.33+2.155*20-0.0433*20^2=20.45 (please try 20.34 if this comes wrong)